Solve for x
x
- + 9x= -9
3

Found 2 solutions by Mike, JimMarshall:
Answer by Mike(39)

1) x + 27x = -27 <---- Multiply through by 3
2) 28x = -27
3) x = -27/28
Hope that helps.
Regards,
-Mike

Answer by JimMarshall(82)

x/3 +9x = -9
==> (x + 27x)/3 = -9 ( LCM of x/3 and 9x/1 = 3 )
==> x + 27x = -27
==> 28x = -27
==> x = -27/28
ANS